OpenStudy (jackiem272):

if f(x) is a continuous function and f(2)=7 and f'(2)=-3, then f(2.01) is approximately

OpenStudy (anonymous):

i guess they want you to linearize the function

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] \[y-7=-3(x-2)\] or \(y=-3(x-2)+7\) replace \(x\) by \(2.1\)

OpenStudy (anonymous):

i.e. \[y=-3(.1)+7\]